A tortoise can run with a speed of 0.1m/s and a hare can run 20 times as fast. In a race, they both start at the same time but the hare stops to rest for 2.0minutes. The tortoise wins by a shell (20cm). (a) how long does the race take? (b) what is the length of the race?
1
Expert's answer
2020-04-30T10:42:14-0400

a)


"d_h=d_t-0.2\\d_t=v_tt\\d_h=v_h(t-120)=20v_t(t-120)\\"

"20(0.1)(t-120)=(0.1)t-0.2"

"t=126\ s"

b)


"d_t=12.6\ m"